Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
bersanetti, pinto - 0:34 Friday 09 April 2021 (51361) Print this report
DRMI recovery: lock with 1f signals & PR alignment in full bandwidth, possible issue with 3f signals

Today we continued the work for the recovery of the DRMI:

  • in order to tune the B2_6MHz_phi0 and B2_56MHz_phi0 demodulation phases we used the PRMI configuration; we needed minimal changes on the 6MHz (final value 1.26) and a quite large change on the 56MHz one; once we managed to lock, clean data is 300 s starting from 16:13:10 UTC;
  • in order to fine tune the 56MHz demodulation phase, we injected noise on PRCL and tried to minimize its coupling to MICH (Figure 1); the results and GPSs are reported in Figure 2 (all measurements lasted 120 s), the optimal demodulation phase was found to be 0.7;
  • then we moved back to the DRMI configuration, which could not be locked immediately; we found the correct gains and then we managed to lock the DRMI at 18:04:13 UTC, although the first locks lasted only for short periods;
  • once the locks were more stable we moved to the PR alignment loops; the standard demodulation phase seemed good enough and we could close the loops both in Drift Control and, starting from 18:10:35 UTC, also in full bandwidth; in this condition we could improve the tuning of the demodulation phases for the QPD signals (both _V and _H changed by -0.2);
  • during the shift the 11.5 Hz noise did not show up in a macroscopic way;
  • at this point the DRMI lock was very stable and repeatable for the rest of the shift, using the Automation to relock it without changes;
  • we moved to some characterization of the DRMI loops, using the UGF monitor lines (Figure 3); we turned them on:
    • MICH @ 31.1 Hz, ampl = 2e-3, at 18:18:00 UTC, UGF ~ 21 Hz;
    • PRCL @ 67.1 Hz, ampl = 1e-4, at 18:19:30 UTC, UGF ~ 52 Hz;
    • SRCL @ 33.3 Hz, ampl = 5e-2, at 18:30:30 UTC, UGF ~ 2.5 Hz;
    • as for the past trials, the SRCL one needed to be unusually high in order to allow a good measurement;
  • then we tried to recover also the lock with the 3f signals; we used the same technique as of #51156; Figures 4 to 6 show the measurements, done using the same UGF lines as before, but with much higher amplitudes in order to see something on the 3f signals; quite strangely, the ratios between 3f and 1f signals resulted to be strongly different with respect to the past;
  • we tried the handoff three times, but all of these were unsuccessfull and resulted in an unlock; Figure 7 shows the last trial, the GPSs are: 20:06:28 UTC, 20:21:04 UTC and 20:30:08 UTC; between one trial and the following one the DRMI relocked automatically without issue, although after the first two unlocks the PR parked itself;
  • we investigated a bit this issue, and we found out that there could be some issue on B2_PD1, which shows signs of possible saturation (Figure 8); the investigation on the 3f signals will continue;
  • all the updated parameters have been stored in the configuration file of the DRMI_LOCK automation node.

At the end of the shift we parked PR and SR and relocked the arm cavities on the IR beam.

Images attached to this report
Comments to this report:
gouaty - 12:40 Friday 09 April 2021 (51369) Print this report
The B2_PD1_RF channel is misbehaving since last tuesday 12h40 utc (Fig.1). At that time we start to observe large values in the RF channel of this photodiode, even though there is no light reaching the PD. A time series showing plenty of loud glitches is shown in Fig.2 at the same moment. The noise spectrum is 3 orders of magnitude larger than what it should be (Fig.3).
This seems to coincide with a reconfiguration of the SIB2 DAQ boxes to solve the timing issue:
Images attached to this comment
gouaty, letendre, masserot - 16:21 Friday 09 April 2021 (51372) Print this report
At 14h02m18-UTC we reconfigured SIB2_DEMOD_B2_PD1_2 in SIB2_dbox_bench.
This solved the problem. The B2 PD1 RF channel is now behaving properly.
Search Help